2 2/3
4 1/4
5 17/20

It's addition

Answer by stanbon(75887)

2 2/3
4 1/4
5 17/20
-----------------------------------
= 11 + (2/3)+(1/4)+(17/20)
-----
= 11 + (40 + 15 + 51)/60
-----
= 11 + (106/60)
---
= 11 + 1 + 46/60
---
= 12 23/30
==============
Cheers,
Stan H.
==============